OpenStudy (anonymous):

If "n" is a positive integer and (n+1)/(2^n)=1/2, then what is the value of "n"? 1 2 3 4 5

OpenStudy (anonymous):

\[\frac{ n+1 }{ 2^n }=\frac{ 1 }{ 2 }\] \[2n+1=2^n\]

OpenStudy (anonymous):

\[2n+2=2^n\]

OpenStudy (anonymous):

 :)